a man and his brother have a combined age of 40 years. 5 years ago the man was 4 times as old as his brother. how old is the man? how old is the brother
Answer by stanbon(75887)

a man and his brother have a combined age of 40 years. 5 years ago the man was 4 times as old as his brother. how old is the man? how old is the brother
---
Equation:
m + b = 40
(m-5) = 4(b-5)
----
Substitute for m and solve for "b":
(40-b-5) = 4b-20
25-b = 4b-20
5b = 45
b = 9 yrs ( brother's age now)
------------------
Solve for "m":
m + b = 40
m + 9 = 40
m = 31 yrs (man's age now)
===================
Cheers,
Stan H.
=================